OpenStudy (anonymous):

Solve: 3(2x - 6) = -4(x - 3) Answer 6 0.6 3 1.8

OpenStudy (tkhunny):

Have you considered the Distributive Property?

OpenStudy (anonymous):

6x-18=-4x+12 now u solve it find the ans.

OpenStudy (anonymous):

do u know multipying bracket term with a no.?

OpenStudy (anonymous):

charlie u should even try to solve it urself

OpenStudy (uri):

Solve the parentheses first 3(2x - 6) = -4(x - 3) 6x-18=-4x+12 Add 18 both sides 6x-18=-4x+12 18 18 ============ 6x=-4x+30 6x+4x=30 10x=30 Divide 30 with 10 x=30/10 x=3 @charlieberzak

OpenStudy (anonymous):

thank you :)